A rectangle is 2 cm longer than it is wide. The diagonal of the rectangle is 10 cm long. Find the perimeter of the rectangle.
Answer by jorel1380(3719)

n2+(n+2)2=(10)2
n2+n2+4n+4=100
2n2+4n-96=0
n2+2n-48=0
(n+8)(n-6)=0
n=6 or -8
The rectangle is 6 cm. by 8 cm.
2(6)+2(8)=P=28 cm
The perimeter is 28 cm..